# RMT-BVQA: Recurrent Memory Transformer-based Blind Video Quality Assessment for Enhanced Video Content

### v2

@@ Abstract
With recent advances in deep learning, numerous algorithms have been developed to enhance video quality, reduce visual artefacts and improve perceptual quality. However, little research has been reported on the quality assessment of enhanced content - the evaluation of enhancement methods is often based on quality metrics that were designed for compression applications. In this paper, we propose a novel blind deep video quality assessment (VQA) method specifically for enhanced video content. It employs a new Recurrent Memory Transformer (RMT) based network architecture to obtain video quality representations, which is optimised through a novel content-quality-aware contrastive learning strategy based on a new database containing 13K training patches with enhanced content. The extracted quality representations are then combined through linear regression to generate video-level quality indices. The proposed method, RMT-BVQA, has been evaluated on the VDPVE (VQA Dataset for Perceptual Video Enhancement) database through a five-fold cross validation. The results show its superior correlation performance when compared to ten existing no-reference quality metrics.
@@ 1 Introduction
Video content is now everywhere! It is by far the largest global Internet bandwidth consumer, with a wide range of applications including consumer video, video conferencing, and gaming. It has been reported that in 2022, each individual in the United Kingdom spent 4.5 hours per day (on average) consuming video content on different platforms [43]. Due to various conditions associated with video capture, editing, and delivery, streamed content often contains a range of visual artefacts, which can affect the quality of a user’s experience. To address this issue, enhancement approaches have been developed, with the aim of reducing visible artefacts and improving overall perceptual quality, e.g. colour transform [16, 34], deblurring [60, 25], deshaking [15], post processing [12, 57], super resolution [9, 56], etc. In particular, more recently, driven by the advances of deep generative models [36, 10, 6], we have seen more effective methods proposed that offer promising enhancement results.
@@
Conventional blind VQA methods [39, 41, 50] are predominantly based on various features extracted in the spatio-temporal or/and frequency domains. Recent advances in this research area favour deep learning-based models employing convolutional neural networks (CNNs) [1, 30] or Vision Transformers (ViTs) [52]. However, these methods tend to exhibit inconsistent performance due to the lack of large and diverse training databases (in particular, for enhanced video content) and inefficient training methodologies. Furthermore, in order to make use of the limited ground-truth quality labels in existing datasets, end-to-end training at the video level is required, but infeasible due to the computation constraints (e.g. memory) with most existing hardware. As a result, many methods [52, 55] resort to sub-sampling videos before evaluating their quality, leading to a significant loss of information. Moreover, it is also noted that most of these blind VQA models were originally optimised and validated on compressed videos that contain artefact types and (spatio-temporal) distributions that differ compared to enhanced content. This can lead to inconsistent metric performance when used to assess enhancement video content.
Figure 1: Illustration of the proposed RMT-BVQA framework.
@@
Inspired by recent works in contrastive learning [37, 59] and the recurrent memory mechanism [2], this paper proposes a blind VQA method, RMT-BVQA (illustrated in Figure 1), based on a new content-quality-aware self-supervised learning methodology and a novel Recurrent Memory Vision Transformer (RMViT) module. This approach first generates a comprehensive representation of the video through dynamically processing global and local information across video frames using the proposed RMViT module. The video representation is then used to predict the final sequence quality index through linear regression. To facilitate the optimisation of the RMViT module through contrastive learning, we have also created a large-scale database containing content generated by various enhancement methods. The primary contributions of this work are summarised as follows.
@@
  • Recurrent Memory Vision Transformer (RMViT) module: To characterise the artefacts exhibiting in the enhanced video content, we designed a new network architecture based on the recurrent memory mechanism [2], which has been employed in language modelling [2], for capturing both short-term temporal dynamics and long-term global information. This also aligns well with the visual persistence characteristic of the human visual system [8]. Moreover, this facilitates end-to-end training at the video level, without sub-sampling the video content, and the recurrence nature allows our model to evaluate videos of variable length. This is the first time when the recurrent memory mechanism is used for the video quality assessment task.
@@
  • Content-quality-aware self-supervised learning: We developed a new self-supervised learning strategy based on a content-quality-aware loss function. This training methodology enables us to optimise the proposed RMViT module based on a large amount of training material without performing expensive and time consuming subjective tests. Here we, for the first time, used a proxy perceptual quality metric to support the quality-aware contrastive learning, inspired by the ranking-based training strategies proposed in [13, 45].
@@
  • A new training dataset: We developed a large and diverse training dataset containing various types of enhanced video content to support the proposed content-quality-aware contrastive learning strategy.
@@
The proposed method has been evaluated on the VQA Dataset for Perceptual Video Enhancement (VDPVE) [14] and achieves superior performance over ten existing no-reference VQA methods based on a five-fold cross validation experiment, with an average Spearman Ranking Order Correlation Coefficient (SRCC) of 0.8209.
@@ 2 Related Work
Objective quality assessment is one of the most important research topics in the field of image processing. It aims to accurately predict the perceptual quality of an input signal given (full reference) or without (no reference) the corresponding reference content. Objective quality metrics play an essential role in comparing different image/video processing methods and supporting algorithm optimisation, e.g. in the rate quality optimisation for compression [42] or as loss functions for learning-based approaches [36]. In the context of video enhancement, since the quality assessment of enhanced content is typically performed when the original reference video is absent, we primarily focus on the no-reference scenario here.
@@
Conventional no-reference quality assessment methods often employ hand-crafted models to observe the input content through spatial, temporal and/or frequency feature extraction. For example, V-BLIINDS [46] employs a spatio-temporal natural scene statistics (NSS) model to quantify motion coherency in video scenes; TLVQM [24] calculates features at various scales from a selection of representative video frames; VIDEVAL [50] predicts video quality by extracting various spatio-temporal artefact features such as motion, jerkiness and blurriness. Other notable contributions include NIQE [41], BRISQUE [39] and V-CORNIA [54]. A more comprehensive review on blind image and video quality metrics can be found in [47].
@@
Learning-based no-reference quality assessment has become more popular recently, inspired by advances in machine learning. Early attempts [23] use various regression models to fit and predict quality indices based on extracted features and conventional quality metrics. More recently, deep neural networks have been utilised for both feature extraction and quality regression to offer improved prediction performance, with important examples including VSFA [30], BVQA [29], SimpleVQA [26], FAST-VQA [52], TB-VQA [53] and SB-VQA [20].
@@
Training methodology. For deep learning-based quality models, it is key to have a large, diverse and representative training database. However, creating such databases is costly since groundtruth quality labels are typically annotated through subjective tests involving human participants. To address this issue, Feng et al. [13] used proxy quality metrics for training content labelling and developed a ranking-inspired training strategy to maintain the reliability of the quality annotation. Moreover, self-supervised learning methods have also been employed which convert the quality labelling into an auxiliary task [37, 38, 59].
@@
Quality assessment for enhanced video content is an underexplored research topic. Previous works typically fine-tune existing blind quality models using enhanced content. A grand challenge [32] was organised in 2023 specifically for enhanced video quality assessment, based on a public training database (VDPVE) [14] which contains enhanced video sequences generated through contrast enhancement, deshaking and deblurring.
@@ 3 Proposed Method
The proposed RMT-BVQA framework is illustrated in Figure 1. It first takes each frame of the input video and transforms it into a one-dimensional embedding using a Spatial Feature Encoder. The extracted embeddings are then sequentially processed by the RMViT (Recurrent Memory Vision Transformer) module to generate the representation of the input video. Finally, a linear ridge regression is employed to output the final sequence quality index. The network architectures employed for each module in this framework, the training database, and the model optimisation strategy are described below in detail.
@@ 3.1 Network architecture
Spatial Feature Encoder. Here we employ a pre-trained ResNet-50 based [18] network which was optimised in a deep image quality model [37] for spatial feature extraction. The network parameters are fixed during our training process for the proposed RMT-BVQA. For each frame, a 2048$\times$1 embedding is extracted.
RMViT module. Considering that artefact types and distributions in the enhanced video content are different from those in compressed content (on which most blind VQA methods are optimised and validated), we designed a new Recurrent Memory Vision Transformer (RMViT) module to effectively capture both the local temporal dynamic and global information within the input video sequence. This is inspired by the recurrent memory mechanism [2] that has been successfully integrated into language models. This mechanism can process and “remember” both local and global information and pass the “memory” between segments within a long sequence using the recurrence structure [2], thus is applicable to videos of any length. As far as we know, this is the first time when the recurrent memory mechanism is applied for the quality assessment task.
As shown in Figure 1, in the first recurrent iteration, the proposed RMViT module takes $N$ frame embeddings ($\mathbf{h}_{1}\ \mathbf{h}_{2}\dots\mathbf{h}_{N}$, corresponding to the first segment in the video) extracted by the Spatial Feature Encoder together with empty memory ($\mathbf{h}_{\mathrm{mem}}^{0}\in\mathbb{R}^{2048\times M}$) as input, where $N$ is a configurable hyperparameter that defines the length of each video segment and $M$ denotes the number of the memory tokens. Specifically, for the first iteration, the input of RMViT is given below:
@@
Regression. The final output of the RMViT module, $\textbf{h}^{v}$, is passed to a linear ridge regressor (the same as in [38]) in order to obtain the sequence level quality index, $Q_{pred}$. Here the regressor is optimised during the cross validation experiment, in which the model parameters of the RMViT are fixed.
@@ 3.2 Training Database Generation
To support the training of the RMViT module, we have collected 149 source videos (with a spatial resolution of 1080p or 720p) from five publicly available datesets including BVI-DVC [35], KoNViD-1k [19], LIVE-VQC [48], Live-Qualcomm [17], and YouTube-UGC [51]. Three primary visual artefacts [^1] The investigation of other types of enhanced content was not the focus of this work, but remains as future work. (as defined in VDPVE [14]) including (i) colour / brightness / contrast degradation, (ii) camera shaking and (iii) blurring, are synthesised (for (iii), through Gaussian filtering) or inherent within the source content (for (i) and (ii)). These videos are then processed using 12 different conventional and learning-based enhancement methods to generate a total of 596 enhanced video sequences. The detailed database generation process is summarised in Table 1.
@@
Table 1: Training database generation.
Each enhanced video and its corresponding reference counterpart are then randomly segmented with a non-overlapping spatio-temporal sliding window to produce training patches with the size of 256 (height)$\times$256 (width)$\times$3 (channel)$\times$72 (temporal length). The reference patches here are only used to generate pseudo-labels for quality classification and are not input into the model. This results in 13,156 (enhanced and reference) patch pairs.
@@ 3.3 Training Strategy
In self-supervised learning, when the model is difficult to optimise directly for its primary task, it is often trained to perform a related pretext task, which can be learnt more effectively. In our case, since assessing the quality of enhanced video content is a challenging task and the corresponding labels are difficult to obtain, inspired by contrastive learning [21, 27], we employ a projector network (with two layers of multilayer perceptron), $g(\cdot)$, to transform the original task into two classification problems focusing on content and quality classification, rather than predicting absolute quality indices.
@@
To enable contrastive learning, a batch ($2B$) of training patches are fed into the network with $B$ randomly selected patches of size 256$\times$256$\times$3$\times$72 and their corresponding down-sampled versions (128$\times$128$\times$3$\times$72). The latter is used here to provide true positive pairs (for the calculation of the quality-aware loss) in contrastive learning as in [37]. The contrastive loss function contains two components, quality-aware and content-aware losses.
@@
Here $P_{i}$ represents the number of patches (positive pairs) in a batch with the same quality interval (the classification process is described below) as patch $i$, or with the similar content to patch $i$ but in different resolutions. $\phi$ stands for the normalised dot product function, $\phi(a,b)=a^{T}b\ /(\|a\|_{2}\|b\|_{2})$, which measures the similarity between $a$ and $b$. $\tau$ is a temperature parameter that is less than 1.
Quality classification. To support quality-aware learning, we employ a proxy quality metric to perform quality classification, inspired by the ranking-based training methodology proposed in RankDVQA [13]. Specifically, we first use VMAF [31] to calculate the quality index for each patch by comparing it with the associated corresponding reference (see subsection 3.2). During the training process, for patch $i$, if another patch $j$ is associated with a VMAF value close to that of patch $i$ (the difference is smaller than a threshold $\mathrm{TH}$),
@@
Content-aware loss captures the content dependent nature of video enhancement methods, providing a different observation aspect in the training process. This has been reported to be effective in the literature [5] for the video quality assessment task. In our case, the positive pair from the content perspective is defined for patches with the same source content.
@@
where $\lambda_{1}$ is a tuning parameter. Here we only consider full-resolution patches when calculating the final contrastive loss.

### v5

# One Cathedral Square, Bristol, BS1 5DD, United Kingdom

@@ Abstract
With recent advances in deep learning, numerous algorithms have been developed to enhance video quality, reduce visual artifacts, and improve perceptual quality. However, little research has been reported on the quality assessment of enhanced content - the evaluation of enhancement methods is often based on quality metrics that were designed for compression applications. In this paper, we propose a novel blind deep video quality assessment (VQA) method specifically for enhanced video content. It employs a new Recurrent Memory Transformer (RMT) based network architecture to obtain video quality representations, which is optimized through a novel content-quality-aware contrastive learning strategy based on a new database containing 13K training patches with enhanced content. The extracted quality representations are then combined through linear regression to generate video-level quality indices. The proposed method, RMT-BVQA  [^1] The code can be downloaded from: https://jasminepp.github.io/RMTBVQA/., has been evaluated on the VDPVE (VQA Dataset for Perceptual Video Enhancement) database through a five-fold cross validation. The results show its superior correlation performance when compared to ten existing no-reference quality metrics.
Keywords:
Video Quality Assessment RMT-BVQA Enhanced Video Content Recurrent Memory Transformer
@@ 1 Introduction
Figure 1: Illustration of the proposed RMT-BVQA framework.
Video content is now everywhere! It is by far the largest global Internet bandwidth consumer, with a wide range of applications including consumer video, video conferencing, and gaming. It has been reported that in 2022, each individual in the United Kingdom spent 4.5 hours per day (on average) consuming video content on different platforms [47]. Due to various conditions associated with video capture, editing, and delivery, streamed content often contains a range of visual artifacts, which can affect the quality of a user’s experience. To address this issue, enhancement approaches have been developed, with the aim of reducing visible artifacts and improving overall perceptual quality, e.g. color transform [17, 38], deblurring [68, 29], deshaking [16], post processing [13, 65], super resolution [10, 63], etc. In particular, more recently, driven by the advances of deep generative models [40, 11, 7], we have seen more effective methods proposed that offer promising enhancement results.
@@
Conventional blind VQA methods [43, 45, 55] are predominantly based on various features extracted in the spatio-temporal or/and frequency domains. Recent advances in this research area favor deep learning-based models employing convolutional neural networks (CNNs) [1, 34] or Vision Transformers (ViTs) [59]. However, these methods tend to exhibit inconsistent performance due to the lack of large and diverse training databases (in particular, for enhanced video content) and inefficient training methodologies. Furthermore, in order to make use of the limited ground-truth quality labels in existing datasets, end-to-end training at the video level is required, but infeasible due to the computation constraints (e.g. memory) with most existing hardware. As a result, many methods [59, 62] resort to subsampling videos before evaluating their quality, leading to a significant loss of information. Moreover, it is also noted that most of these blind VQA models were originally optimized and validated on compressed videos that contain artifact types and (spatio-temporal) distributions that differ compared to enhanced content. This can lead to inconsistent metric performance when used to assess enhanced video content.
@@
Inspired by recent works in contrastive learning [41, 67] and the recurrent memory mechanism [3], we proposes a blind VQA method, RMT-BVQA (illustrated in Fig. 1), based on a new content-quality-aware self-supervised learning methodology and a novel Recurrent Memory Vision Transformer (RMViT) module. This approach first generates a comprehensive video representation through dynamically processing global and local information across video frames using the proposed RMViT module. The video representation is then used to predict the final sequence quality score through linear regression. To facilitate the optimization of the RMViT module through contrastive learning, we have also created a large-scale database containing content generated by various enhancement methods. The primary contributions of this work are summarized as follows.
  • 1)
@@
    Recurrent Memory Vision Transformer (RMViT) module: To characterize the artifacts exhibiting in the enhanced video content, we designed a new network architecture based on the recurrent memory mechanism [3], which has been employed in language modeling [3], for capturing both short-term temporal dynamics and long-term global information. This also aligns well with the visual persistence characteristic of the human visual system [9]. Moreover, this facilitates end-to-end training at the video level, without sub-sampling the video content, and the recurrence nature allows our model to evaluate videos of variable length. This is the first time when the Recurrent Memory Vision Transformer [^2] We acknowledge that other recurrent mechanisms, such as LSTMs, have been employed for VQA in the literature [19]. is used for the video quality assessment task.
  • 2)
@@
    Content-quality-aware self-supervised learning: We developed a new self-supervised learning strategy based on a content-quality-aware loss function. This training methodology enables us to optimize the proposed RMViT module based on a large amount of training material without performing expensive and time consuming subjective tests. Here we, for the first time, used a proxy perceptual quality metric to support the quality-aware contrastive learning for VQA, inspired by the ranking-based training strategies [14, 49].
  • 3)
@@
    A new training dataset: We developed a large and diverse training dataset containing various types of enhanced video content to support the proposed content-quality-aware contrastive learning strategy. This training dataset is made available for future research.
@@
The proposed method has been evaluated on the VQA Dataset for Perceptual Video Enhancement (VDPVE) [15] and achieves superior performance over ten existing no-reference VQA methods based on five-fold cross validations, with an average Spearman Rank-order Correlation Coefficient (SRCC) of 0.8209.
@@ 2 Related Work
Objective quality assessment is one of the most important research topics in the field of image processing. It aims to accurately predict the perceptual quality of an input signal given (full reference [58, 35, 64]) or without (no reference [45, 26, 59]) the corresponding reference content [^3] Another type of quality assessment methods does exist, denoted reduced-reference models, which only employ partial information from the reference.. Objective quality metrics play an essential role in comparing different image/video processing methods and supporting algorithm optimization, e.g. in the rate quality optimization for compression [46] or as loss functions for learning-based approaches [40]. In the context of video enhancement, since the quality assessment of enhanced content is typically performed when the original reference video is absent, we primarily focus on the no-reference scenario here.
@@
Conventional no-reference quality assessment methods often employ hand-crafted models to observe the input content through spatial, temporal and/or frequency feature extraction. For example, V-BLIINDS [51] employs a spatio-temporal natural scene statistics (NSS) model to quantify motion coherency in video scenes; TLVQM [28] calculates features at various scales from a selection of representative video frames; VIDEVAL [55] predicts video quality by extracting various spatio-temporal artifact features such as motion, jerkiness, and blurriness. Other notable contributions include NIQE [45], BRISQUE [43], and V-CORNIA [61]. A more comprehensive review of blind image and video quality metrics can be found in [52].
@@
Learning-based no-reference quality assessment has become more popular recently, inspired by advances in machine learning. Early attempts [26, 51, 28, 12] use various regression models to fit and predict quality indices based on extracted features and conventional quality metrics. More recently, deep neural networks have been utilized for both feature extraction and quality regression to offer improved prediction performance, with important examples including BVQA [33], SimpleVQA [30], FAST-VQA [59], TB-VQA [60] and SB-VQA [23].
@@
Training methodology For deep learning-based quality models, it is key to have a large, diverse, and representative training database. However, creating such databases is costly, since ground-truth quality labels are typically annotated through subjective tests involving human participants. To address this issue, Feng et al. [14] used proxy quality metrics for labeling training material and developed a ranking-inspired training strategy to maintain the reliability of quality annotation. Moreover, self-supervised learning methods have also been employed that convert quality labeling into an auxiliary task [41, 42, 67].
@@
Quality assessment for enhanced video content is an underexplored research topic. Previous works typically fine-tune existing blind quality models using enhanced content. A grand challenge [36] was organized in 2023 specifically for enhanced video quality assessment, based on a public training database (VDPVE) [15], which contains enhanced video sequences generated through contrast enhancement, deshaking and deblurring. However, it should be noted that only the training set in the VDPVE database is publicly available, while the test dataset used in the grand challenge has not been released.
@@ 3 Proposed Method
The proposed RMT-BVQA framework is illustrated in Fig. 1. It first takes each frame of the input video and transforms it into a one-dimensional embedding using a Spatial Feature Encoder. The extracted embeddings are then sequentially processed by the RMViT (Recurrent Memory Vision Transformer) module to generate the representation of the input video. Finally, a linear ridge regression is employed to output the final predicted sequence quality score. The network architectures employed for each module in this framework, the training database, and the model optimization strategy are described below in detail.
@@ 3.1 Network architecture
Spatial Feature Encoder. Here we employ a pre-trained ResNet-50 based [20] network which was optimized in a deep image quality model [41] for spatial feature extraction. The network parameters are fixed during our training process for the proposed RMT-BVQA. For each frame, a 2048$\times$1 embedding is extracted.
RMViT module. Considering that artifact types and distributions in enhanced video content are different from those in compressed content (on which most blind VQA methods are optimized and validated), we designed a new Recurrent Memory Vision Transformer (RMViT) module to effectively capture both local temporal dynamic and global information within the input video sequence. This is inspired by the recurrent memory mechanism [3] that has been successfully integrated into language models. This mechanism can process and “remember" both local and global information and pass the “memory” between segments within a long sequence using the recurrence structure [3], thus is applicable to videos of any length. To our knowledge, this is the first time the recurrent memory mechanism has been applied to the quality assessment task.
As shown in Fig. 1, in the first recurrent iteration, the proposed RMViT module takes $N$ frame embeddings ($\mathbf{h}_{1}\ \mathbf{h}_{2}\dots\mathbf{h}_{N}$, corresponding to the first segment in the video) extracted by the Spatial Feature Encoder together with empty memory ($\mathbf{h}_{\mathrm{mem}}^{0}\in\mathbb{R}^{2048\times M}$) as input, where $N$ is a configurable hyperparameter that defines the length of each video segment and $M$ denotes the number of memory tokens. Specifically, for the first iteration, the input of RMViT is given below:
@@
Regression. The final output of the RMViT module, $\textbf{h}^{v}$, is passed to a linear ridge regressor (the same as in [42]) in order to obtain the predicted sequence level quality score, $Q_{pred}$. Here, the regressor is optimized during the cross-validation experiment, in which the model parameters of the RMViT are fixed.
@@ 3.2 Training Database Generation
Table 1: Training database generation.
@@
To support the training of the RMViT module, we have collected 149 source videos (with a spatial resolution of 1080p or 720p) from five publicly available datasets including BVI-DVC [39], KoNViD-1k [22], LIVE-VQC [53], Live-Qualcomm [18], and YouTube-UGC [57]. Three primary visual artefacts [^4] The investigation of other types of enhanced content was not the focus of this work, but remains as future work. (as defined in VDPVE [15]) including (i) color / brightness / contrast degradation, (ii) camera shaking, and (iii) blurring are synthesized (for (iii), through Gaussian filtering) or inherent within the source content (for (i) and (ii)). These videos are then processed using 12 different conventional and learning-based enhancement methods to generate a total of 596 enhanced video sequences. The detailed database generation process is summarized in Tab. 1.
Each enhanced video and its corresponding reference counterpart are then randomly segmented with a non-overlapping spatio-temporal sliding window to produce training patches with the size of 256 (height) $\times$ 256 (width) $\times$ 3 (channel) $\times$ 72 (temporal length). The reference patches here are only used to generate pseudo-labels for quality classification and are not input into the model. This results in 13,156 (enhanced and reference) patch pairs. Fig. 2 shows example frames of the training sequences generated for training the proposed model.
Figure 2: Example frames of the training sequences. The first row shows example frames corresponding to the videos generated using four different enhancement methods for color, contrast and brightness enhancement. The second row presents the examples from the content generated by four deshaking methods. The third row shows those produced by four deblurring methods.
@@ 3.3 Training Strategy
In self-supervised learning, when the model is difficult to optimize directly for its primary task, it is often trained to perform a related pretext task, which can be learned more effectively. In our case, since assessing the quality of enhanced video content is a challenging task and the corresponding labels are difficult to obtain, inspired by contrastive learning [24, 31], we employ a projector network (with two layers of multilayer perceptron), $g(\cdot)$, to transform the original task into two classification problems focusing on content and quality classification, rather than predicting absolute quality indices.
@@
To enable contrastive learning, a batch ($2B$) of training patches are fed into the network with $B$ randomly selected patches of size 256$\times$256$\times$3$\times$72 and their corresponding down-sampled versions (128$\times$128$\times$3$\times$72). The latter is used here to provide true positive pairs (for the calculation of the quality-aware loss) in contrastive learning, as in [41]. The contrastive loss function contains two components, quality-aware and content-aware losses.
@@
Here $P_{i}$ represents the number of patches (positive pairs) in a batch with the same quality interval (the classification process is described below) as patch $i$, or with the similar content to patch $i$ but in different resolutions. $\phi$ stands for the normalized dot product function, $\phi(a,b)=a^{T}b\ /(\|a\|_{2}\|b\|_{2})$, which measures the similarity between $a$ and $b$. $\tau$ is a temperature parameter that is less than 1.
Quality classification. To support quality-aware learning, we employ a proxy quality metric to perform quality classification, inspired by the ranking-based training methodology proposed in RankDVQA [14]. Specifically, we first use VMAF [35] to calculate the quality score for each patch by comparing it with the associated corresponding reference (see Sec. 3.2). During the training process, for patch $i$, if another patch $j$ is associated with a VMAF value close to that of patch $i$ (the difference is smaller than a threshold $\mathrm{TH}$),
@@
Content-aware loss captures the content-dependent nature of video enhancement methods, providing a different observation aspect in the training process. This has been reported to be effective in the literature [6] for the video quality assessment task. In our case, the positive pair from the content perspective is defined for patches with the same source content.
@@
where $\lambda_{1}$ is a tuning parameter. Here, we only consider full-resolution patches when calculating the final contrastive loss.